# Kapranov degrees

### v3

@@ 4. Upper bound
Let us first recall the general observation that a Kapranov degree $\int_{\overline{M}_{0,n}}X_{S_{1},i_{1}}\dotsm X_{S_{n-3},i_{n-3}}$ is zero when $\bigcup_{i=1}^{n-3}S_{i}\neq[n]$ because in that case $X_{S_{1},i_{1}}\dotsm X_{S_{n-3},i_{n-3}}$ is the pullback of a product of $n-3$ divisors on $\overline{M}_{0,\bigcup_{i}S_{i}}$, which has dimension at most $n-4$.
@@
in the summation, for a partition $[n]=P\sqcup Q$ with $P\ni 1$, is nonzero only if $P=\{1,b\}$ for some $b\neq 1$ since $1\notin T_{j}$ for all $j>1$. When $P=\{1,b\}$ for some $b\neq 1$, we find that the term evaluates to 1 as follows: $\int_{\overline{M}_{0,P\cup\star}}\prod_{j=2}^{n-3}\operatorname{res}_{P}(T_{j},j)=1$ because $\operatorname{res}_{P}(T_{j},j)=1$ for all $j>1$ and $\overline{M}_{0,P\cup\star}$ is a point, while $\int_{\overline{M}_{0,Q\cup\star}}\prod_{j=2}^{n-3}\operatorname{res}_{Q}(T_{j},j)=1$ by Lemma 4.2 because $\operatorname{res}_{Q}(T_{j},j)=X_{\{j,\star,n-1,n\}}$ if $b=n-2$ (and likewise if $b=n-1$ or $n$), and otherwise $\operatorname{res}_{Q}(T_{j},j)=X_{\{j,n-2,n-1,n\}}$ for $j\neq b$ and $X_{\{\star,n-2,n-1,n\}}$ for $j=b$.
Now, if $1\in S\setminus i$, then applying Theorem B with $a=1$, we find $\int_{\overline{M}_{0,n}}X_{S\setminus 1,i}X_{T_{2}}\dotsm X_{T_{n-3}}=0$ by the general observation, and all the terms in the summation over the appropriate partitions $[n]=P\sqcup Q$ vanish except when $P=\{1,i\}$, which contributes 1.
@@
Note that subsets $S_{1},\dotsc,S_{n-3}$ of $[n-2]$ satisfy the dragon marriage condition if and only if $S_{1}\cup\{n-1,n\},\dotsc,S_{n-3}\cup\{n-1,n\}$ satisfies the Cerberus condition.
@@
    For subsets $S_{1},\dotsc,S_{n-3}$ of $[n-2]$, the Kapranov degree
@@
For a partition $[n]=P\sqcup Q$ with $|P|\geq 2$, $n-2\in P$, and $\{n-1,n\}\subseteq Q$, define a collection $\mathcal{T}|_{P}$ of subsets of $P\setminus\{n-2\}$, and a collection $\mathcal{T}|_{Q}$ of pairs $(T,j)$ where $T\subseteq Q\setminus\{n-1,n\}$ and $j\in\{n-2,n-1,n\}$, by
$\mathcal{T}|_{P}$ =$=\{T_{i}\cap P:j_{i}=n-2\text{ and }T_{i}\cap P\neq\varnothing\}\quad\text{and}$
$\mathcal{T}|_{Q}$ =$=\{(T_{i}\cap Q,j_{i}):j_{i}=n-2\text{ and }T_{i}\cap P=\varnothing\}\cup\{(T_{i}\cap Q,j_{i}):j_{i}\in\{n-1,n\}\}.$
@@
$\operatorname{gl}_{P|Q}^{*}Y_{T_{i},j_{i}}=\begin{cases}\pi_{P}^{*}X_{T_{i}\cap P\cup\{n-2,\star\},n-2}&\text{if }j_{i}=n-2\text{ and }T_{i}\cap P\not=\varnothing\\ \pi_{Q}^{*}X_{T_{i}\cup\{n-1,n,\star\},\star}&\text{if }j_{i}=n-2\text{ and }T_{i}\cap P=\varnothing\\ \pi_{Q}^{*}X_{T_{i}\cap Q\cup\{n-1,n,\star\},j_{i}}&\text{otherwise, i.e.\ }j_{i}\in\{n-1,n\},\end{cases}$
@@
where for the second factor about $\overline{M}_{0,Q\cup\{n-2\}}$, we relabeled $\star$ by $n-2$. The lemma now follows from Proposition 5.1(2). ∎
@@
[…]
@@
for all $U_{1}\supseteq T_{1}\cup a$, $U_{2}\supseteq T_{2},\dotsc,U_{n-3}\supseteq T_{n-3}$. This contradicts the fact we have equality when $U_{i}=[n-3]$ for all $i$.
@@
where the sum is over all partitions $[n]=P\sqcup Q$ with $\{a,n-2\}\subseteq P$ and $T_{1}\cup\{n-1,n\}\subseteq Q$ such that the subsets of $P\setminus(n-2)$ in $\mathcal{T}|_{P}$ satisfy the dragon marriage condition. Given such a partition $[n]=P\sqcup Q$ and a $3$-transversal $t$ of $\mathcal{T}|_{Q}$, extend $t$ to a $3$-transversal $\tilde{t}$ of $(T_{1}\cup a,n-2),\dotsc,(T_{n-3},j_{n-3})$ by setting $\tilde{t}(i)=n-2$ if $i\in P\setminus(n-2)$, and $\tilde{t}(i)=t(i)$ if $i\in Q\setminus\{n-1,n\}$. This is indeed a $3$-transversal: if $m$ is a bijection $\{i:(T_{i},j_{i})\in\mathcal{T}|_{Q}\}\to Q\setminus\{n-1,n\}$ that witnesses the 3-transversal $t$, then we can extend $m$ to $\tilde{m}\colon[n-3]\to[n-3]$ by setting $\tilde{m}(1)=a$ and using Proposition 5.1(1) to extend $\tilde{m}$ to the remaining indices $\{i:T_{i}\in\mathcal{T}|_{P}\}$.
In order to prove the claimed bound, we need to check that the $3$-transversal obtained is new, and that no such new $3$-transversal can be obtained from two different partitions $[n]=P\sqcup Q$. For verifying this, we note the following observation.
@@
Lastly, suppose now that there is some other partition $[n]=P^{\prime}\sqcup Q^{\prime}$ which also gives rise to $\tilde{t}$ via the above construction. Then, by the observation, a bijection $\tilde{m}$ realizing $\tilde{t}$ must restrict to give a bijection $\{i:j_{i}=n-2\text{ and }T_{i}\cap P^{\prime}=\varnothing\}\to\{b\in Q^{\prime}\setminus\{n-1,n\}:\tilde{t}(b)=n-2\}$. As $Q^{\prime}\not=Q$, we thus have $\{i:j_{i}=n-2\text{ and }T_{i}\cap P^{\prime}=\varnothing\}\not=\{i:j_{i}=n-2\text{ and }T_{i}\cap P=\varnothing\}$. Hence, by switching $(P|Q)$ with $(P^{\prime}|Q^{\prime})$ if necessary, we may assume that
$W=\{i:j_{i}=n-2,\enspace T_{i}\cap P=\varnothing,\text{ and }T_{i}\cap P^{\prime}\not=\varnothing\}$
@@
When $k=0$ or $n-3$, these mixed volumes were computed in [Pos09, Theorem 5.1], and when $n-2\in T_{j}$ for all $j$, they were computed in [EFLS, Theorem B], which is generalized by Theorem D. It can be deduced from [HK12, BES] that when all of the simplices are of the form $\Delta_{T}$ for some $T$ or $\nabla_{[n-2]}$, then this mixed volume can be computed in terms of the characteristic polynomial of a certain cotransversal matroid. It then follows from [CPV] that computing Kapranov degrees is #P-hard.

### v4

@@ 4. Upper bound
Let us first recall the general observation that a Kapranov degree $\int_{\overline{M}_{0,n}}X_{S_{1},i_{1}}\dotsm X_{S_{n-3},i_{n-3}}$ is zero when $\bigcup_{i=1}^{n-3}S_{i}\neq[n]$ because in that case $X_{S_{1},i_{1}}\dotsm X_{S_{n-3},i_{n-3}}$ is the pullback of a product of n3n-3 divisors on $\overline{M}_{0,\bigcup_{i}S_{i}}$, which has dimension at most n4n-4.
@@
in the summation, for a partition $n]=P\sqcup Q$ with $P\ni 1$, is nonzero only if $P=\{1,b\}$ for some $b\neq 1$ since $1\notin T_{j}$ for all $j>1$. When $P=\{1,b\}$ for some $b\neq 1$, we find that the term evaluates to 1 as follows: $\int_{\overline{M}_{0,P\cup\star}}\prod_{j=2}^{n-3}\operatorname{res}_{P}(T_{j},j)=1$ because $\operatorname{res}_{P}(T_{j},j)=1$ for all $j>1$ and $\overline{M}_{0,P\cup\star}$ is a point, while $\int_{\overline{M}_{0,Q\cup\star}}\prod_{j=2}^{n-3}\operatorname{res}_{Q}(T_{j},j)=1$ by Lemma 4.2 because $\operatorname{res}_{Q}(T_{j},j)=X_{\{j,\star,n-1,n\}}$ if b=n2b=n-2 (and likewise if b=n1b=n-1 or $n$), and otherwise $\operatorname{res}_{Q}(T_{j},j)=X_{\{j,n-2,n-1,n\}}$ for $j\neq b$ and $X_{\{\star,n-2,n-1,n\}}$ for $j=b$.
Now, if $1\in S\setminus i$, then applying Theorem B with $a=1$, we find $\int_{\overline{M}_{0,n}}X_{S\setminus 1,i}X_{T_{2}}\dotsm X_{T_{n-3}}=0$ by the general observation, and all the terms in the summation over the appropriate partitions $n]=P\sqcup Q$ vanish except when $P=\{1,i\}$, which contributes 1.
@@ 5. $3$-transversals
Note that subsets $S_{1},\dotsc,S_{n-3}$ of [n2][n-2] satisfy the dragon marriage condition if and only if $S_{1}\cup\{n-1,n\},\dotsc,S_{n-3}\cup\{n-1,n\}$ satisfies the Cerberus condition.
@@
    For subsets $S_{1},\dotsc,S_{n-3}$ of [n2][n-2], the Kapranov degree
@@
For a partition $n]=P\sqcup Q$ with $|P|\geq 2$, $n-2\in P$, and $\{n-1,n\}\subseteq Q$, define a collection $\mathcal{T}|_{P}$ of subsets of $P\setminus\{n-2\}$, and a collection $\mathcal{T}|_{Q}$ of pairs $(T,j)$ where $T\subseteq Q\setminus\{n-1,n\}$ and $j\in\{n-2,n-1,n\}$, by
[…]
$\mathcal{T}|_{Q}$ $=\{(T_{i}\cap Q,j_{i}):j_{i}=n-2\text{ and }T_{i}\cap P=\varnothing\}\cup\{(T_{i}\cap Q,j_{i}):j_{i}\in\{n-1,n\}\}.$
@@
[…]
@@
where for the second factor about $\overline{M}_{0,Q\cup\{n-2\}}$, we relabeled $\star$ by n2n-2. The lemma now follows from Proposition 5.1(2). ∎
@@
$\int_{\overline{M}_{0,n}}Y_{T_{1}\cup a,n-2}\dotsb Y_{T_{n-3},j_{n-3}}-\int_{\overline{M}_{0,n}}Y_{T_{1},n-2}\dotsb Y_{T_{n-3},j_{n-3}}\leq|\{\text{new 3-transversals}\}|,$
@@
for all $U_{1}\supseteq T_{1}\cup a$, $U_{2}\supseteq T_{2},\dotsc,U_{n-3}\supseteq T_{n-3}$. This contradicts the fact that we have equality when $U_{i}=[n-3]$ for all $i$.
@@
where the sum is over all partitions $n]=P\sqcup Q$ with $\{a,n-2\}\subseteq P$ and $T_{1}\cup\{n-1,n\}\subseteq Q$ such that the subsets of $P\setminus(n-2)$ in $\mathcal{T}|_{P}$ satisfy the dragon marriage condition. Given such a partition $n]=P\sqcup Q$ and a $3$-transversal $t$ of $\mathcal{T}|_{Q}$, extend $t$ to a $3$-transversal $\tilde{t}$ of $(T_{1}\cup a,n-2),\dotsc,(T_{n-3},j_{n-3})$ by setting $\tilde{t}(i)=n-2$ if $i\in P\setminus(n-2)$, and $\tilde{t}(i)=t(i)$ if $i\in Q\setminus\{n-1,n\}$. This is indeed a $3$-transversal: if $m$ is a bijection $\{i:(T_{i},j_{i})\in\mathcal{T}|_{Q}\}\to Q\setminus\{n-1,n\}$ that witnesses the 3-transversal $t$, then we can extend $m$ to $\tilde{m}\colon[n-3]\to[n-3]$ by setting $\tilde{m}(1)=a$ and using Proposition 5.1(1) to extend $\tilde{m}$ to the remaining indices $\{i:T_{i}\in\mathcal{T}|_{P}\}$.
In order to prove the claimed bound, we need to check that the $3$-transversal obtained is new, and that no such new $3$-transversal can be obtained from two different partitions $n]=P\sqcup Q$. For verifying this, we note the following observation.
@@
Lastly, suppose now that there is some other partition $n]=P^{\prime}\sqcup Q^{\prime}$ which also gives rise to $\tilde{t}$ via the above construction. Then, by the observation, a bijection $\tilde{m}$ realizing $\tilde{t}$ must restrict to give a bijection $\{i:j_{i}=n-2\text{ and }T_{i}\cap P^{\prime}=\varnothing\}\to\{b\in Q^{\prime}\setminus\{n-1,n\}:\tilde{t}(b)=n-2\}$. As $Q^{\prime}\not=Q$, we thus have $\{i:j_{i}=n-2\text{ and }T_{i}\cap P^{\prime}=\varnothing\}\not=\{i:j_{i}=n-2\text{ and }T_{i}\cap P=\varnothing\}$. Hence, by switching $(P|Q)$ with $(P^{\prime}|Q^{\prime})$ if necessary, we may assume that
$W=\{i:j_{i}=n-2,\enspace T_{i}\cap P=\varnothing,\text{ and }T_{i}\cap P^{\prime}\not$$\varnothing\}$
@@
When $k=0$ or n3n-3, these mixed volumes were computed in [Pos09, Theorem 5.1], and when $n-2\in T_{j}$ for all $j$, they were computed in [EFLS, Theorem B], which is generalized by Theorem D. It can be deduced from [HK12, BES] that when all of the simplices are of the form $\Delta_{T}$ for some $T$ or $\nabla_{[n-2]}$, then this mixed volume can be computed in terms of the characteristic polynomial of a certain cotransversal matroid. It then follows from [CPV] that computing Kapranov degrees is #P-hard.